# Written by AAAI Press Staff1

### v1

@@ 3.1 Learning Policy $\pi$ as Markov Decision Process
The above reward mechanism provides flexibility to our framework to be versatile to objectives of different nature, that may or may not be differentiable. This is an advantage over existing neural method (nazi2019gap) where having a differentiable form of the partitioning objective is a pre-requisite.
3.2 Policy Loss Computation and Parameter Update
Our objective is to learn parameters of our policy network in such a way that actions that lead to an overall improvement of the partitioning objective are favored more over others. Towards this, we use REINFORCE gradient estimator (williams1992simple) to optimize the parameters of our policy network. Specifically, we wish to maximize the reward obtained for the horizon of length $T$ with discounted rewards $D^{t}$. Towards this end, we define a reward function $J(\pi_{\theta})$ as:
@@ 3.2 Policy Loss Computation and Parameter Update
3.3 Training and Inference
For a given set of training graphs, we optimize the parameters of the policy network $\pi_{\theta}$ for $T$ steps. Subsequently, given a new graph and partition objective, we adapt the policy parameters $\pi_{\theta}$ using Eq. 16. Note that the trajectory length $T$ is not kept very large to avoid the long-horizon problem.
During inference, we compute the initial node embeddings, obtain initial partitioning and then run the forward pass of our policy to improve the partitioning objective over time.
@@ 3.4 Time complexity
$\mathcal{O}\big{(}(\alpha\times\beta)+(|\mathcal{E}|+k)\times T^{\prime}\big{)}$. Here $\alpha$ is number of anchor nodes, $\beta$ is number of random walk iterations, $k$ is the number of partitions and $T^{\prime}$ is the number of iterations during inference. Typically $\alpha$, $\beta$ and $k$ are $<<|\mathcal{V}|$ and $T^{\prime}=o(|\mathcal{V}|)$. Further, for sparse graphs $\mathcal{E}=\mathcal{O}(|\mathcal{V}|)$. Hence time complexity of NeuroCUT is $o(|\mathcal{V}|^{2})$. In comparison to spectral clustering, which is a common graph partitioning algorithm having time complexity of $\mathcal{O}(|\mathcal{V}|^{3})$, the proposed technique’s complexity is much lower. This can also be attributed to the fact that NeuroCUT exploits the sparse nature of graphs unlike spectral clustering that performs matrix operations of $\mathcal{O}(|\mathcal{V}|^{3})$. For detailed derivation please see appendix  C.
@@
Figure 2: Results of all methods on five datasets through radar plots. The color shade represents identity of each method. The legend in the first subplot is common for all plots. All values (for each partitioning objective) are normalized by dividing by the worst (maximum) cut value obtained on that dataset. The smaller the value, i.e, closer to the center, better is the performance. Further, when two polygons are overlapping, but not fully contained within each other, in general, the smaller polygon tends to indicate a better holistic performance across all objectives. Fig (a) represents cut values averaged across five datasets. Fig (b-f) depict results on each of the five datasets. The neural state-of-the-art GAP method produces the most inferior performance among considered baselines. Further in several cases it generates ”NaN” values (error). Our method NeuroCUT  either outperforms the baselines or produces the second best results while being close to the best one in all cases. Moreover, NeuroCUT  is the most robust and does not produce the worst cut values. The neural methods are trained and tested with the same number partitions ($k=5$). Unnormalized results for each dataset in table format are presented in Table  7 in appendix.
4 Experiments
  • Efficacy and robustness: NeuroCUT consistently produces the best or the second best results in terms of the average performance over all objective functions. This establishes the robustness of NeuroCUT.
  • Inductivity: As one of the major strengths, unlike the state-of-the-art neural model GAP (nazi2019gap), NeuroCUT is inductive on the number of partitions and, consequently, generalizes well over unseen number of partitions.
4.1 Experimental Setup
Datasets:
We use five datasets for our experiments among which one is synthetic. They are as follows:
  • Cora and Citeseer (sen2008collective): These are citation networks. In these, a $k$-way cut could represent the partitioning of publications into $k$ clusters while minimizing the citations across the clusters. Each cluster might correspond to a distinct research field, shedding light on interdisciplinary connections or isolating specialized domains.
  • Facebook (leskovec2012learning): In social network analysis, a $k$-way cut can delineate communities of users with minimal inter-community connections. This partitioning unveils hidden social circles and provides insights into how different groups interact within the larger network. The datasets have multiple connected components, we work with the largest connected component to make sure we can evaluate proper cuts.
  • Harbin (li2019learning): In this road network, a $k$-way cut could entail partitioning road segments into $k$ subsets to optimize traffic flow between them. This partitioning aids in urban planning by identifying key intersections and routes that minimize congestion.
  • Stochastic Block Models (SBM) (abbe2017community): This synthetic network is generated by SBM where $k$-way cuts reveal the underlying community structure. This helps to validate our method by comparing the identified partitions with the ground truth communities.
Baselines:
We compare our proposed method NeuroCUT with the state-of-the-art neural method called GAP (nazi2019gap). It allows balanced normalized cuts with an end-to-end framework with a differentiable loss function which is a continuous relaxation version of normalized cut. We also use two popular non-neural baselines: hMetis (karypis1997multilevel) and spectral clustering (ng2001spectral).
Other settings:
We run all our experiments on an Ubuntu 20.04 system running on Intel Xeon 6248 processor with 96 cores and 1 NVIDIA A100 GPU with 40GB memory for our experiments. We used GraphSage(hamilton2017inductive) as our GNN with number of layers $L=2$, learning rate as $0.0001$, hidden size = $32$. We used Adam optimizer for training the parameters of our policy network $\pi_{\theta}$. For computing discounted reward in RL, we use discount factor $\gamma=0.99$. We set the length of trajectory $T$ during training as $2$. At time step $t$, the rewards are computed from time $t$ to $t+T$ and parameters of the policy $\pi_{\theta}$ are updated. The default number of anchor nodes for computing positional embeddings is set to $35$.
4.2 Results on Transductive Setting
We compare our proposed method NeuroCUT against all the baselines where the neural models are trained and tested with the same number of partitions ($k=5$). Fig. 2 presents the results for all datasets and all methods. The results are depicted using spider plots where the smaller the value i.e closer to the center, better is the performance. Further, in the plots, if two polygons are overlapping, but not fully contained within each other, in general the smaller polygon tends to indicate a better holistic performance across all objectives (heckert2002handbook). NeuroCUT outperforms the baselines or produces the next to optimal results. Moreover, NeuroCUT is the most robust among all the methods and does not produce the worst cut values in any of the settings. Overall, the performance of NeuroCUT is the most consistent among all the methods. Unnormalized results for each dataset in table format are presented in Table  7 in appendix.
NeuroCUT incorporates causality during inference and this leads to the robust performance observed in fig. 2. Specifically, as the architecture is auto-regressive, it takes into account the current state before moving ahead as opposed to a single shot pass in GAP (nazi2019gap). We observe that the non-neural methods fail to generalize well across all objective functions since they do not learn the base heuristic as a function of the objective and the dataset. Finally, we note that in SBMs the task is easier as there are already $5$ clusters which aligns well with the the desired number of partitions, which is also $5$.
4.3 Results on Inductive Setting
Dataset Method $normalised$ $sparsest$ $balanced$ $kmin$
Cora hMetis 0.772 0.772 3.177 0.097
Spectral 0.723 1.503 1.845 0.008
NeuroCUT 0.591 2.451 0.981 0.022
CiteSeer hMetis 0.415 0.415 1.554 0.056
Spectral 0.328 1.094 0.792 0.003
NeuroCUT 0.363 1.060 0.457 0.031
SBMs hMetis 3.259 3.260 65.955 0.393
Spectral 2.495 2.513 50.805 0.242
NeuroCUT 3.241 5.505 45.602 0.038
Table 2: Inductive results based on the number of partitions of all methods on all six datasets. The best ones are shown by green and the second best ones are in gray. Lower values are better. Our method NeuroCUT either outperforms the baselines or produces the second best results while being close to the best one in ten out of twelve settings. NeuroCUT is trained on different partition sizes ($k=5$ and $k=10$) and the test data involves an unseen partition size ($k=8$). GAP is not inductive in terms of number of partitions and thus the results are not present here.
We show that NeuroCUT generalizes to unknown number of partitions. We compare it against only the non-neural baselines since GAP does not generalize to unseen number of partitions. Here, NeuroCUT is trained on different partition sizes ($k=5$ and $k=10$) and the test data involves an unseen partition size ($k=8$). Table 2 shows the results. Our method NeuroCUT either outperforms the baselines or produces the second best results while being close to the best one in most of the settings. Note that while the non-neural methods need to be re-run for the unseen partitions, NeuroCUT only runs forward pass to produce the results on an unseen partition size.
4.4 Generalization to Topological changes in graphs
In this section, we analyze the performance of NeuroCUT to topological changes in the graph. Specifically, once a model is trained, we aim to evaluate its performance when the topology of the graph changes which is a common scenario in dynamic networks such as social networks where connections are formed/deleted over time.
Towards this first we train NeuroCUT on a given graph. Next, we add/remove $x\%$ percentage of edges in the graph. Then we predict the cuts on the perturbed graph. To predict cuts for the perturbed graph, our model NeuroCUT only needs to just run a forward pass over it.
Tables  3,4,5,6 show the results on two datasets, namely, Cora and Facebook for different partitioning objectives at different degrees of perturbations. In almost all settings, our method NeuroCUT outperforms the baselines. The only neural baseline i.e GAP had to be retrained on perturbed graph and still produced NaN values(indicated by -) in most of the cases. These results show that our proposed method NeuroCUT is more robust in comparison to existing neural methods at different levels of perturbations. Additionally, in several cases it outperforms non-neural methods which were directly optimized on perturbed graphs.
[…]
Cora hMetis No 0.351 0.355 0.399 0.435 0.506
Spectral No 0.346 0.442 0.479 0.424 0.490
GAP Yes 0.512 0.500 0.483 0.808 0.664
NeuroCUT Yes 0.321 0.307 0.364 0.428 0.569
Facebook hMetis No 1.112 1.113 1.126 1.126 1.131
Spectral No 2.034 1.676 1.918 1.92 1.679
GAP Yes - - 0.482 - -
NeuroCUT Yes 0.487 0.459 0.463 0.494 0.864
Table 3: Perturbing $x\%$ Edges - Normalised Cuts
[…]
Cora hMetis No 1.442 1.464 1.649 1.819 2.169
Spectral No 0.795 1.058 1.224 0.957 1.452
GAP Yes 2.029 1.978 1.915 3.167 2.816
NeuroCUT Yes 1.015 1.317 1.479 1.546 2.053
Facebook hMetis No 51.669 52.937 52.110 52.962 52.236
Spectral No 5.625 4.722 5.591 5.591 4.867
GAP Yes - - 11.886 - -
NeuroCUT Yes 3.528 3.516 3.745 4.060 5.298
Table 4: Perturbing $x\%$ Edges - Sparsest Cuts
[…]
Cora hMetis No 0.352 0.355 0.399 0.435 0.506
Spectral No 1.089 1.223 0.800 1.207 0.841
GAP Yes 0.536 0.529 0.515 0.870 0.681
NeuroCUT Yes 0.526 0.544 0.613 0.651 0.836
Facebook hMetis No 1.112 1.113 1.126 1.127 1.131
Spectral No 2.818 2.459 2.697 2.697 2.456
GAP Yes - - 0.506 - -
NeuroCUT Yes 0.831 0.793 0.800 0.855 0.921
Table 5: Perturbing $x\%$ Edges - Balanced Cuts
[…]
Cora hMetis No 0.071 0.072 0.080 0.087 0.101
Spectral No 0.002 0.001 0.054 0.001 0.064
GAP Yes 0.099 0.097 0.094 0.114 0.126
NeuroCUT Yes 0.010 0.013 0.014 0.018 0.021
Facebook hMetis No 0.199 0.203 0.201 0.203 0.201
Spectral No 0.000 0.000 0.000 0.000 0.000
GAP Yes 0.024 0.026 0.044 0.026 0.027
NeuroCUT Yes 0.000 0.000 0.000 0.000 0.000
Table 6: Perturbing $x\%$ Edges - KMin Cuts
[…]
Figure 3: Results on the initial and the final cut values at $k=5$. It shows that our neural model NeuroCUT (Final) performs more accurate node and partition selection to optimize the objective function. Subsequently, there is a significant difference between the initial and final cut values.
4.5 Ablation Studies
Initial vs Final Cut Values. NeuroCUT takes a warm start by partitioning nodes based on clustering over their Lipschitz embeddings, which are subsequently fine-tuned auto-regressively through the reinforcement learning based framework. How much does the partitioning objective improve since the initial clustering? Figure 3 sheds light on this question. Specifically, it shows the difference between initial cut value due after clustering and the final cut value from the partitions produced by our method. We note that there is a significant difference between the initial and final cut values, which essentially shows the effectiveness of NeuroCUT.
[…]
Figure 4: NeuroCUT results with heuristics for node selection in Phase 1 for $k=5$. Our heuristic finds better cuts than a random selection algorithm for two objective functions: Normalized Cut and KMin Cut.
Impact of Node Selection Procedures. Figure 4 shows how our proposed heuristic for node selection helps in finding the near-optimal cuts for any given objective function. We compare and show that a simpler node selection where we select all the nodes one by one in a random order produces substantially inferior results. Furthermore, the difference lies not only in final cut-values but also we require much less perturbations.
In App. D of appendix, we present additional results on variation of parameters of NeuroCUT.

### v2

@@ 3.3 Reward, Training & Inference
The above reward mechanism provides flexibility to our framework to be versatile to objectives of different nature, that may or may not be differentiable. This is an advantage over existing neural methods (nazi2019gap; tsitsulin2023graph) where having a differentiable form of the partitioning objective is a pre-requisite.
Policy Loss Computation and Parameter Update. Our objective is to learn parameters of our policy network in such a way that actions that lead to an overall improvement of the partitioning objective are favored more over others. Towards this, we use REINFORCE gradient estimator (williams1992simple) to optimize the parameters of our policy network. Specifically, we wish to maximize the reward obtained for the horizon of length $T$ with discounted rewards $D^{t}$. Towards this end, we define a reward function $J(\pi_{\theta})$ as:
@@
Training and Inference. For a given graph, we optimize the parameters of the policy network $\pi_{\theta}$ for $T$ steps. Note that the trajectory length $T$ is not kept very large to avoid the long-horizon problem. During inference, we compute the initial node embeddings, obtain initial partitioning and then run the forward pass of our policy to improve the partitioning objective over time.
@@ 3.4 Time complexity
$\mathcal{O}\big{(}(\alpha\times\beta)+(|\mathcal{E}|+k)\times T^{\prime}\big{)}$. Here $\alpha$ is the number of anchor nodes, $\beta$ is number of random walk iterations, $k$ is the number of partitions and $T^{\prime}$ is the number of iterations during inference. Typically $\alpha$, $\beta$ and $k$ are $<<|\mathcal{V}|$ and $T^{\prime}=o(|\mathcal{V}|)$. Further, for sparse graphs $\mathcal{E}=\mathcal{O}(|\mathcal{V}|)$. Hence time complexity of NeuroCUT is $o(|\mathcal{V}|^{2})$. For detailed derivation please see appendix  A.4.
4 Experiments
  • Efficacy and robustness: NeuroCUT produces the best results over diverse partitioning objective functions. This establishes the robustness of NeuroCUT.
  • Inductivity: As one of the major strengths, unlike existing neural models such as GAP (nazi2019gap), DMon (tsitsulin2023graph) and MinCutPool (bianchi2020spectral), our proposed method NeuroCUT is inductive on the number of partitions and consequently can generalize to unseen number of partitions.
4.1 Experimental Setup
Datasets:
We use four real datasets for our experiments. They are described below and their statistics are present in the appendix (please see Table 7).
  • Cora and Citeseer (sen2008collective): These are citation networks where nodes correspond to individual papers and edges represent citations between papers. The node features are extracted using a bag-of-words approach applied to paper abstracts.
  • Harbin (li2019learning): This is a road network extracted from Harbin city, China. The nodes correspond to road intersections and node features represent latitude and longitude of a road intersection.
  • Actor (platonov2023critical): This dataset is based upon Wikipedia data where each node in the graph corresponds to an actor, and the edge between two nodes denotes co-occurrence on the same Wikipedia page. Node features correspond to keywords in Wikipedia pages.
Partitioning objectives:
We evaluate our method on a diverse set of four partitioning objectives described in Section  2, namely Normalized Cut, Balanced Cut, k-MinCut, and Sparsest Cut. In addition to evaluating on diverse objectives, we also choose a diverse number of partitions for evaluation, specifically, $k=2,\;5\;\text{and}\;10$.
Baselines:
We compare our proposed method with both neural as well as non-neural methods.
Neural baselines: We compare with DMon (tsitsulin2023graph), GAP (nazi2019gap), MinCutPool  (bianchi2020spectral) and Ortho (tsitsulin2023graph). DMon is the state-of-the-art neural attributed-graph clustering method. GAP is optimized for balanced normalized cuts with an end-to-end framework with a differentiable loss function which is a continuous relaxation version of normalized cut. MinCutPool optimizes for normalized cut and uses an additional orthogonality regularizer. Ortho is the orthogonality regularizer described in DMon and MinCutPool.
Non-neural baselines: Following the settings of DMon (tsitsulin2023graph), we compare with K-means clustering applied on raw node features.
Other settings:
We run all our experiments on an Ubuntu 20.04 system running on Intel Xeon 6248 processor with 96 cores and 1 NVIDIA A100 GPU with 40GB memory for our experiments. For NeuroCUT we used GraphSage(hamilton2017inductive) as our GNN with number of layers $L=2$, learning rate as $0.0001$, hidden size = $32$. We used Adam optimizer for training the parameters of our policy network $\pi_{\theta}$. For computing discounted reward in RL, we use discount factor $\gamma=0.99$. We set the length of trajectory $T$ during training as $2$. At time step $t$, the rewards are computed from time $t$ to $t+T$ and parameters of the policy $\pi_{\theta}$ are updated. The default number of anchor nodes for computing positional embeddings is set to $35$.
Dataset Method $k=2$ $k=5$ $k=10$
Cora Kmeans 0.65 3.26 7.44
MinCutPool 0.12 0.61 1.65
DMon 0.57 3.07 7.40
Ortho 0.80 1.88 4.06
GAP 0.10 0.68 -
NeuroCUT 0.02 0.33 0.92
CiteSeer Kmeans 0.30 2.35 5.21
MinCutPool 0.10 0.38 1.04
DMon 0.33 2.71 6.84
Ortho 0.28 1.51 3.25
GAP 0.12 - -
NeuroCUT 0.02 0.20 0.44
Harbin Kmeans 0.56 2.34 5.40
MinCutPool - - -
DMon 0.98 3.25 -
Ortho - - -
GAP 0.25 - -
NeuroCUT 0.01 0.07 0.28
Actor Kmeans 0.99 4.00 8.98
MinCutPool 0.55 1.97 4.73
DMon 0.77 3.46 8.08
Ortho 1.05 3.98 8.90
GAP 0.20 - -
NeuroCUT 0.17 0.99 4.66
Table 2: Results on Normalized Cut. Our model NeuroCUT produces the best (lower is better) performance across all datasets and the number of partitions $k$.
Dataset Method $k=2$ $k=5$ $k=10$
Cora Kmeans 3.41 14.90 32.52
MinCutPool 0.52 2.44 6.68
DMon 0.45 1.89 5.82
Ortho 2.73 7.06 15.70
GAP 0.41 2.80 -
NeuroCUT 0.13 1.46 3.03
CiteSeer Kmeans 1.53 13.70 22.20
MinCutPool 0.31 1.32 3.62
DMon 0.35 1.21 3.62
Ortho 0.88 4.15 10.60
GAP 0.60 - -
NeuroCUT 0.11 0.49 1.19
Harbin Kmeans 1.91 7.44 17.03
MinCutPool - - -
DMon 2.01 - -
Ortho - - -
GAP 1.56 - -
NeuroCUT 0.06 0.23 0.82
Actor Kmeans 5.43 19.40 40.34
MinCutPool 2.44 8.51 19.81
DMon 1.71 9.50 21.01
Ortho 3.42 16.55 40.4
GAP 1.35 - -
NeuroCUT 0.65 2.04 2.88
Table 3: Results on Sparsest Cut. Our model NeuroCUT produces the best (lower is better) performance across all datasets and number of partitions $k$.
4.2 Results on Transductive Setting
In the transductive setting, we compare our proposed method NeuroCUT against the mentioned baselines, where the neural models are trained and tested with the same number of partitions. Tables 2-5 present the results for different partitioning objectives across all datasets and methods. For the objectives under consideration, a smaller value depicts better performance. NeuroCUT demonstrates superior performance compared to both neural and non-neural baselines across four distinct partitioning objectives, highlighting its robustness. Further, we would also like to point out that in many cases, existing baselines produce “nan” values which are represented by “-” in the result tables.
Our method NeuroCUT incorporates dependency during inference and this leads to robust performance. Specifically, as the architecture is auto-regressive, it takes into account the current state before moving ahead as opposed to a single shot pass in methods such as GAP (nazi2019gap). Further, unlike other methods, NeuroCUT also incorporates positional information in the form of Lipschitz embedding to better contextualize global node positional information. We also observe that the non-neural method K-means fails to perform on all objective functions since its objective is not aligned with the main objectives under consideration.
Dataset Method $k=2$ $k=5$ $k=10$
Cora Kmeans 0.32 0.34 0.61
MinCutPool 0.06 0.12 0.17
DMon 0.05 0.09 0.14
Ortho 0.33 0.34 0.38
GAP 0.05 0.10 0.11
NeuroCUT $\sim$ 0.0 $\sim$ 0.0 0.06
CiteSeer Kmeans 0.12 0.27 0.43
MinCutPool 0.04 0.08 0.10
DMon 0.05 0.07 0.12
Ortho 0.13 0.47 0.30
GAP 0.05 0.08 0.10
NeuroCUT 0.01 0.02 0.03
Harbin Kmeans 0.28 0.46 0.54
MinCutPool 0.0 0.0 0.0
DMon 0.10 0.01 $\sim$ 0.0
Ortho 0.0 0.0 0.0
GAP 0.01 0.01 0.01
NeuroCUT $\sim$ 0.0 0.01 0.03
Actor Kmeans 0.48 0.54 0.80
MinCutPool 0.25 0.35 0.42
DMon 0.17 0.39 0.44
Ortho 0.35 0.65 0.83
GAP 0.09 0.12 0.26
NeuroCUT $\sim$ 0.0 $\sim$ 0.0 $\sim$ 0.0
Table 4: Results on k-MinCut. In most of the cases, our model NeuroCUT produces the best (lower is better) performance across all datasets and the number of partitions $k$. Values less than $10^{-2}$ are approximated to 0.
Dataset Method $k=2$ $k=5$ $k=10$
Cora Kmeans 0.68 3.90 7.44
MinCutPool 0.13 0.60 1.62
DMon 0.11 0.48 1.47
Ortho 0.80 1.89 4.06
GAP 0.10 0.74 -
NeuroCUT 0.45 0.64 1.08
CiteSeer Kmeans 0.42 2.64 5.30
MinCutPool 0.09 0.38 1.04
DMon 0.10 0.37 1.1
Ortho 0.28 1.51 3.32
GAP 0.23 - -
NeuroCUT 0.07 0.24 0.60
Harbin Kmeans 0.56 2.37 5.41
MinCutPool - - -
DMon 1.30 - -
Ortho - - -
GAP 0.72 - -
NeuroCUT 0.21 0.11 0.27
Actor Kmeans 1.00 4.06 9.08
MinCutPool 0.55 1.96 4.71
DMon 0.34 2.01 4.80
Ortho 1.05 3.90 8.91
GAP 0.25 - -
NeuroCUT 0.59 1.69 4.42
Table 5: Results on Balanced Cut. In most cases, our model NeuroCUT produces the best (lower is better) performance across all datasets and number of partitions $k$.
4.3 Results on Inductivity to Partition Count
Dataset Method Normalized Sparsest k-MinCut Balanced
Cora Kmeans 7.44 32.52 0.61 7.53
NeuroCUT-T 0.92 3.03 0.06 1.08
NeuroCUT-I 1.18 5.04 0.02 2.03
CiteSeer Kmeans 5.21 22.20 0.43 5.30
NeuroCUT-T 0.44 1.19 0.03 0.60
NeuroCUT-I 0.63 1.62 0.03 0.62
Harbin Kmeans 5.40 17.03 0.54 5.41
NeuroCUT-T 0.28 0.82 0.03 0.27
NeuroCUT-I 0.27 0.87 0.03 0.29
Table 6: Inductivity to unseen partition count. Here we set the target number of partitions $k=10$. I stands for inductive and T for transductive setting. In this table, the transductive version of NeuroCUT trained on $k=10$ serves as a point of reference when assessing the quality of the inductive version of NeuroCUT.
@@ 4.3 Results on Inductivity to Partition Count
Figure 2: Results on the initial warm start and the final cut values obtained by NeuroCUT at $k{=}10$. It shows that our neural model NeuroCUT (Final) performs more accurate node and partition selection to optimize the objective function. Subsequently, there is a significant difference between the initial and final cut values.
As detailed in Section 3.2, the decoupling of parameter size and the number of partitions allows NeuroCUT to generalize effectively to an unseen number of partitions. In this section, we empirically analyze the generalization performance of NeuroCUT to an unknown number of partitions. We compare it with the non-neural baseline K-means, as the neural methods like GAP, DMon, Ortho, and MinCutPool cannot be employed to infer on an unseen number of partitions. In Table 6, we present the results on inductivity. Specifically, we trained NeuroCUT on different partition sizes ($k=5$ and $k=8$) and tested it on an unseen partition size $k=10$. The inductive version, referred to as NeuroCUT-I [^2] For clarity purposes, in this experiment we use NeuroCUT-T to refer to the transductive setting whose results are presented in Sec.  4.2, obtains high-quality results on different datasets. Note that while the non-neural method such as K-means needs to be re-run for the unseen partitions, NeuroCUT only needs to perform forward pass to produce the results on an unseen partition size.
4.4 Ablation Studies
Initial Warm Start vs Final Cut Values. NeuroCUT takes a warm start by partitioning nodes based on clustering over their raw node features and Lipschitz embeddings, which are subsequently fine-tuned auto-regressively through the phase 1 and phase 2 of NeuroCUT. In this section, we measure, how much the partitioning objective has improved since the initial clustering? Figure 2 sheds light on this question. Specifically, it shows the difference between the initial cut value after clustering and the final cut value from the partitions produced by our method. We note that there is a significant difference between the initial and final cut values, which essentially shows the effectiveness of NeuroCUT.
Figure 3: Node Selection in Phase 1 with Heuristic vs Random: Relative % improvement (gain) in cut values obtained by NeuroCUT when using node selection heuristic against random node selection for $k=5$. In most cases, our heuristic finds significantly better cuts than a random node selection procedure.